# Atomizer: Generalizing to new modalities by breaking satellite images down to a set of scalars

### v1

@@ 5 Results
This section presents experimental results comparing Atomizer with baseline models (Perceiver jaegle2021perceiver, ViT dosovitskiy2020image, ResNet he2016deep, and ScaleMAE reed2023scale) across different settings, resolutions, and input sizes.
@@
Table 3 presents model performance across input sizes from 30$\times$30 to 120$\times$120 pixels. Atomizer’s performance increases with input size, from 37.98% AP at 30$\times$30 pixels to 48.66% AP at 120×120 pixels (a 28.1% relative improvement). At all input sizes, Atomizer outperforms all models. ResNet shows non-monotonic behavior, with performance peaking at 60$\times$60 pixels (31.30% AP) before declining at 120$\times$120 pixels (29.50% AP). ViT and ScaleMAE exhibit more consistent scaling patterns but with lower overall performance than Atomizer. The Perceiver model maintains constant performance (15.16% AP) regardless of input size. This suggests that despite its token-based design, Perceiver does not effectively utilize the additional spatial context in larger images. These results demonstrate the advantage of Atomizer’s approach to encoding spatial positions within each token.
@@
Model 30$\times$30 px 60$\times$60 px 90$\times$90 px 120$\times$120 px

### v2

@@ 5 Results
This section presents experimental results comparing Atomizer with baseline models (Perceiver jaegle2021perceiver (8), ViT dosovitskiy2020image (10), ResNet he2016deep (17), and ScaleMAE reed2023scale (6)) across different settings, resolutions, and input sizes.
@@
Table 3 presents model performance across input sizes from $\times30$ to $\times120$ pixels. Atomizer’s performance increases with input size, from 37.98% AP at $\times30$ pixels to 48.66% AP at 120×120 pixels (a 28.1% relative improvement). At all input sizes, Atomizer outperforms all models. ResNet shows non-monotonic behavior, with performance peaking at $\times60$ pixels (31.30% AP) before declining at $\times120$ pixels (29.50% AP). ViT and ScaleMAE exhibit more consistent scaling patterns but with lower overall performance than Atomizer. The Perceiver model maintains constant performance (15.16% AP) regardless of input size. This suggests that despite its token-based design, Perceiver does not effectively utilize the additional spatial context in larger images. These results demonstrate the advantage of Atomizer’s approach to encoding spatial positions within each token.
@@
Model $\times30$ px $\times60$ px $\times90$ px $\times120$ px